# Rome was Not Built in a Single Step: Hierarchical Prompting for LLM-based Chip Design

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: MLCAD; Sept. 9-11, 2024; Snowbird, UT
@@ 1. Introduction
In this paper, we develop and evaluate hierarchical prompting techniques to facilitate automated generation of modular HDL code. We explore hierarchical Verilog generation in two major modalities, each occurring in the real-world. In the human-driven mode, the prompt contains a human-proposed hierarchy that the LLM must extract and implement, as well as iterative compiler feedback from unit tests for each submodule. In the more challenging purely generative mode, the LLM gets only a basic (non-hierarchical) prompt and therefore must make its own design decisions to implement the target module. We implement an 8-stage pipeline to automate these techniques in a generalizable fashion. This allows an LLM to closely emulate human HDL development practices.
@@ 3.1. Sources of Hierarchy
3.2. Hierarchical Generation Pipeline

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024conference: MLCAD; Sept. 9-11, 2024; Snowbird, UTisbn: 979-8-4007-0699-8/24/09
[…]
@@ 1. Introduction
In this paper, we develop and evaluate hierarchical prompting techniques to facilitate automated generation of modular HDL code. We explore hierarchical Verilog generation in two major modalities, each occurring in the real-world. In the human-driven mode, the prompt contains a human-proposed hierarchy that the LLM must extract and implement, as well as iterative compiler feedback from unit tests for each submodule. In the more challenging purely generative mode, the LLM gets only a basic (non-hierarchical) prompt and therefore must make its own design decisions to implement the target module. We implement an 8-stage pipeline to automate these techniques in a generalizable fashion, which we refer to as Recurrent Optimization via Machine Editing (ROME). This allows an LLM to closely emulate human HDL development practices.
@@ 3.1. Sources of Hierarchy
3.2. ROME Hierarchical Generation Pipeline
@@ 7. Conclusion
Appendix A Additional Figures
Top Modules Submodules
64-to-1 Multiplexer
2-to-1 mux
4-to-1 mux
8-to-1 mux
16-to-1 mux
32-to-1 mux
64-to-1 mux
5-to-32 Decoder
2-to-4 decoder
3-to-8 decoder
5-to-32 decoer
32-bit Barrel Shifter
8-bit Barrel Shifter
Rotation Control
32-bit Barrel Shifter
4x4 Systolic Array
Processor Element
Control Logic
Top-level 4x4 Array
8-bit UART
Baud Rate Generator
Receiver
Transmitter
State definitions
Top-level UART
128-bit AES Block Cipher
S-box
Key Memory
Encipher Block
Decipher Block
Control Logic
Helper Functions(e.g.:
inverse shift rows,
mix columns, etc.)
Table 5. Top-level modules and corresponding submodules of the hierarchical benchmarks.
Figure 5. Perseveration-like behavior in GPT-3.5 when asked for rare syntax.
Figure 6. Elaborated Design Schematic of LLM-Generated MIPS Processor. Zoom in for submodule information.
Figure 7. Sample waveform of running a series of instructions. Our first two instructions test the load and store functions, moving values between the registers and data memory. We then load two values from data memory, 12,006 and 13,663, into the registers R4 and R5 respectively. We then test the ADD function to sum them to 25669, or 6445 in hex. This intentionally echoes the value of the instruction for easy confirmation on the waveform. We then load two more values, 38,776 and 1,104 to test the OR instruction, to once again echo the instruction value of 9778 hex. We then test the SUBI instruction by subtracting 35,498 from 38,776 to get 3,278, or 0CCE hex. We then store all of our outputs thus far into data memory at various addresses to confirm successful saving. More thorough testing omitted for brevity. Though the LLM implements opcodes that are notably different from the standard MIPS ISA, all instructions are present and functional. Full instruction series: LW R7, 8(R7); SW R8, 9(R8); LW R4, 1(R4); LW R5, 3(R5); ADD R6, R4, R5; LW R7, 5(R7); LW R8, 6(R8); OR R9, R7, R8; SUBI R10, R9, 35498; SW R6, 9(R0); SW R9, 10(R0); SW R10, 11(R0);
Appendix B Artifact Evaluation
All relevant artifacts, including scripts, are publicly available at:
https://github.com/ajn313/ROME-LLM
They are additionally publicly archived at:
https://zenodo.org/records/13323449
https://figshare.com/projects/ROME-LLM/214771
We provide an example Python notebook to be used in Google Colab that includes all necessary elements for convenience. All that is required is an OpenAI API key with GPT-4 usage enabled. We present a novel algorithm here in the form of our automated design pipeline. It utilizes both an error handling feedback loop in addition to a hierarchical feedback loop in order to build on simpler module to make more complex circuits & structures. The benchmark we utilized is the novel Hierarchical ROME LLM benchmark described within the paper. This includes a number of modules to be tested. We include unit tests to validate the performance of our tool.
All simulations are run using the open-source iVerilog tool which will be downloaded by the Colab notebook and will run via script calls in the Colab environment. No other compiler/simulator is required. All software dependencies are handled by the Colab notebook, but a list of dependencies will additionally be available on the GitHub. The notebook will default to using GPT-4 as the model, though this can easily be changed to another OpenAI model. No download is necessary as it will run remotely through the OpenAI API. Instructions will be included on how to instead use an open-source model, including our own CL-Verilog model, though it will be easiest to test using GPT. Additionally, CL-Verilog will be released on Hugging Face.
Running all benchmarks may take numerous hours, but we include a subset of examples that should be less time consuming while still displaying performance. Measurements can include pass@k, time to completion, or binary pass/fail. These values are affected by the stochasticity of LLMs and may not match up perfectly to our own observations, for better or worse. Results should be reproducible via our Colab notebook, though there is likely to be notable variance between runs, as our pass@k measurements show. The final output should be a functional version of the target module which passes testbench simulation.
The testbenches can be downloaded from GitHub and uploaded to Colab, or cloned directly from the repo. No more than a few megabytes of local disk space should be required if downloading is preferred, as the testbenches are on average around 2kb. One option for quickly opening our notebook is githubtocolab.com, which will open the notebook directly in Google Colab from GitHub. Instructions are on the tool’s site.
MIT license is used for code where applicable. All artifacts will be archived on Zenodo and FigShare and publicly available on GitHub at the above links.
Zenodo DOI:
https://doi.org/10.5281/zenodo.13323449